OpenStudy (anonymous):

can someone help me please im in connexus and im doing the quadratic functions and equations ill give medal for awnsers please help

OpenStudy (anonymous):

dude

OpenStudy (anonymous):

is ur teacher ms.wholwend?

OpenStudy (anonymous):

what grade

OpenStudy (anonymous):

Are u taking the OGT's?

OpenStudy (anonymous):

i go to connexus academy ohio

OpenStudy (anonymous):

Dude i am 12 in 6th grade -_-

OpenStudy (anonymous):

im in algebra 1 and my teachers mr eblan and im indiana connections

OpenStudy (anonymous):

oh cool who r u?

OpenStudy (anonymous):

U fan me i fan u

OpenStudy (anonymous):

i really need to get this done soon or im in trouble so can any body help me

OpenStudy (anonymous):

We cant help u if u dont give us da question :P

OpenStudy (anonymous):

how many real number solutions are there in this quation -8x^2-8x-2=0

OpenStudy (anonymous):

it is either 1 or 2 hold on i just posted it i got recorses :P

OpenStudy (anonymous):

lol ok man

OpenStudy (anonymous):

fan me and best response me plz u fan me i fan u

OpenStudy (anonymous):

ill fan u after u help me :)

OpenStudy (anonymous):

Okay :P

OpenStudy (anonymous):

Is it short answer?

OpenStudy (anonymous):

here

OpenStudy (anonymous):

ok then write this : Δ=64-4*16=0 x=-8/16=-1/2=-0.5

OpenStudy (anonymous):

here it is the formula

OpenStudy (anonymous):

ok then write this : Δ=64-4*16=0 x=-8/16=-1/2=-0.5 now can u fan me

OpenStudy (anonymous):

YAY